Determine the stopping distances for an automobile going a constant initial speed of 93km/h and human reaction time of 0.30 s, for an acceleration a=−3.3m/s2.


1
Expert's answer
2021-06-30T10:15:08-0400

From equation of motion

"v_f^2-v_i^2=2a\Delta x\\\ \\\Delta x=\dfrac{v_f^2-v_i^2}{2a}"


Stopping distance "d=vt+\Delta x"


Here, "v_i=93\ km/h=93(5/18)\ m/s= 26\ m/s"


"\Delta x=\dfrac{v_f^2-v_i^2}{2a}=\dfrac{0^2-26^2}{-2\times 3.3}=102.42\ m"


and Stopping Distance "d=vt+Δx =23(0.30)+102.42=109.32\ m"